### v4

@@ Appendix B LSTM architecture
where $\bm{i}_{t}$, $\bm{f}_{t}$, $\bm{g}_{t}$ and $\bm{o}_{t}$ are the input, forget, cell and output gates, respectively. Each of these gates have their learnable weights for both the inputs $\bm{x}_{t}$ (represented by the matrix $\bm{W})$ and hidden states $\bm{h}_{t}$ (represented by the matrix $\bm{U}$, as well as learnable biases (represented by the vector $\bm{b}$). Observe that these learnable parameters don’t depend on time – even if the result of each gate may depend on time with $\bm{x}_{t}$ and $\bm{h}_{t}$, they use the same learnable parameters, which are applied recursively. The input, forget and output gates are enclosed by the sigmoid function $\sigma$ and the cell gate is enclosed by the hyperbolic tangent function $\tanh$. Furthermore, in the expression of cell state $\bm{c}_{t}$ and the hidden state $\bm{h}_{t}$, $\odot$ represents an element-wise multiplication. Figure 9 schematizes these equations inside the LSTM cell.

### v5

@@ Appendix B LSTM architecture
where $\bm{i}_{t}$, $\bm{f}_{t}$, $\bm{g}_{t}$ and $\bm{o}_{t}$ are the input, forget, cell and output gates, respectively. Each of these gates have their learnable weights for both the inputs $\bm{x}_{t}$ (represented by the matrix $\bm{W})$ and hidden states $\bm{h}_{t}$ (represented by the matrix $\bm{U}$), as well as learnable biases (represented by the vector $\bm{b}$). Observe that these learnable parameters don’t depend on time – even if the result of each gate may depend on time with $\bm{x}_{t}$ and $\bm{h}_{t}$, they use the same learnable parameters, which are applied recursively. The input, forget and output gates are enclosed by the sigmoid function $\sigma$ and the cell gate is enclosed by the hyperbolic tangent function $\tanh$. Furthermore, in the expression of cell state $\bm{c}_{t}$ and the hidden state $\bm{h}_{t}$, $\odot$ represents an element-wise multiplication. Figure 9 schematizes these equations inside the LSTM cell.